^ X+5 =8
FIND X
I NEED HELP ASAP

Answer by queenofit(120)

If you ment x^2 + 5 = 8 I might be able to help.
first get the numbers and the variables on their "team" sides.
x^2= 8-5
x^2= 3
take the square root of x and 3, remember when you take the square root of a number the answer could either be negative or positive so you will have 2 answers
x= the positive or negative square root of 3